OpenStudy (anonymous):

A person travelling 6kmph reaches his office 15 min late.Had he travelled at 8kmph, he would have been 25 minutes early. find the distance

OpenStudy (anonymous):

6(t+15)=d 8(t-25)=d

OpenStudy (anonymous):

dude kmph and and minutes ... so need to convert ... plz solve

OpenStudy (anonymous):

t+1/4 = d/6 t-5/12 = d/8

OpenStudy (anonymous):

now i assume u can solve linear eqns

OpenStudy (anonymous):

Let the distance be x Speed =. ( distance / time ) Distance = speed * time 6t + 90 = 8t -200 2t = 290 T =145 t - 25 = 120 = 2 hrs 120 + 15 = 135 135/60 * 6 = d D = 13.5 km

OpenStudy (anonymous):

answer is 16 km

OpenStudy (anonymous):

solve plz

OpenStudy (anonymous):

??

OpenStudy (dumbcow):

use him1618 equations and solve for d using either substitution or elimination t = d/6 - 1/4 -> (d/6 - 1/4) -5/12 = d/8

OpenStudy (dumbcow):

-1/4 - 5/12 = d/8 - d/6 -3/12 -5/12 = 6d/48 - 8d/48 -8/12 = -2d/48 -2/3 = -d/24 48/3 = d 16 = d

OpenStudy (anonymous):

ya thanks........ :)